OpenStudy (anonymous):

What is the discriminant of 9x2 + 2 = 10x? –356 –172 28 72

OpenStudy (anonymous):

28?

OpenStudy (anonymous):

9x^2 + 2 -10x = 0 ax^2 + bx + c = 0 9x^2 -10x + 2 = 0 sqrt(b^2 -4ac) (100 - 4 (9)(2) (100-72) 28

OpenStudy (anonymous):

great move on